OpenStudy (anonymous):

how do find the derivative of r/(sqrt of r^2 +4)

OpenStudy (anonymous):

use the derivative for division rule: f(x)/g(x) = f'(x)g(x)-f(x)g'(x)/g(x)^2

OpenStudy (anonymous):

i tried that but i keep getting the wrong answer

OpenStudy (anonymous):

= (sqrt of r^2 +4) - r(1/2)(2r)(r^2+4)^-1/2/r^2+4

OpenStudy (anonymous):

=(r^2+4 - r^2) /(r^2+4)^3/2

OpenStudy (anonymous):

=4 /(r^2+4)^3/2

OpenStudy (anonymous):

o i see my mistake i put for the second part (r^2+4)^-3/2

OpenStudy (anonymous):

the final answer is =4 /(r^2+4)^3/2

OpenStudy (anonymous):

could you became my fan?

OpenStudy (anonymous):

im still confused on the part (sqrt of r^2 +4) - r(1/2)(2r)(r^2+4)^-1/2/r^2+4

OpenStudy (anonymous):

how did u get (r^2+4)^-1/2

OpenStudy (anonymous):

i kept getting it to the power of -3/2

OpenStudy (anonymous):

ok... when we derivative a function we always do this for example: the derivative of x^1/2 = 1/2(x)^(1/2-1) = 1/2(x)^-1/2

OpenStudy (anonymous):

apply that for you example: the derivative of (r^2+4)^1/2 = (1/2)(2r)(r^2+4)^(1/2-1) = (r)(r^2+4)^(-1/2)

OpenStudy (anonymous):

do you get it?

OpenStudy (anonymous):

ooooooooooooo

OpenStudy (anonymous):

hahaha ok i get it!! THANK YOU!!!

OpenStudy (anonymous):

your welcome!